How do I solve the following inequality using interval notation? 1/x-3 <= 1/x-5
Answer by stanbon(75887)

How do I solve the following inequality using interval notation?
1/(x-3) <= 1/(x-5)
------
x cannot be 3 and it cannot be 5.
------
Solve the equality.
1/(x-3) = 1/(x-5)
---
x-5 = x-3
No solution: So no solution for the equality part of the problem.
-----------------
Draw a number line and plot x = 3 and x = 5.
---
Test values in the resulting intervals to find inequality solutions.
---
1/(x-3)<1/(x-5)
--
If x = 0 you get 1/-3 < 1/-5: true, so solution in (-inf,3)
If x = 4 you get 1/1 < 1/-1: false
If x = 10 you get 1/7 < 1/5: true, so solutions in (5,+inf)
------
Cheers,
Stan H.